twix01:

A candy bar box is in the shape of a triangular prism. The volume of the box is 1,200 cubic centimeters. A triangular prism is shown with base of triangle labeled 10 cm, sides of the triangles labeled 13 cm, and length of the box equal to 20 cm. Part A: What is the height of the box? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make the sides of the candy box? Explain how you got your answer. (5 points)

twix01:

please help

twix01:

1 attachment
twix01:

@Ultrilliam

twix01:

@JustSaiyan

twix01:

@ThisGirlPretty

twix01:

@Ultrilliam

twix01:

pls help

twix01:

@eatsalad

Shadow:

The easiest way to solve for the volume would be to use the triangle as the base. Then 20cm would be the height.

Shadow:

Proven by: \[A _{T} = \frac{ 13 \times 10 }{ 2 } = 65\] \[V = 65 \times 20 = 1300cm^3\]

Shadow:

So I would use 20cm as the height of the box.

twix01:

but the volume is 1200 not 1300

Shadow:

That doesn't make sense. The height of the box is 20cm. If you find the area of the triangle, then multiply it by 20, you could get the volume ._.

twix01:

but i already have the volume, i just need the height

twix01:

and part b

Shadow:

Lets see if salad has something c;

eatsalad:

A V=1/2*Base area*H x=height 1/2*10*x*20=1200 solve x=12

twix01:

wow, got anything for part b???

Shadow:

Doesn't make any sense though when they give us 13.

Shadow:

Must be a typo

Shadow:

In such a case I would have solved it the same way salad. You're good.

eatsalad:

b 13+10*20=213

eatsalad:

Or 13+13*20

twix01:

wow, your really good

twix01:

wait, which one???

eatsalad:

Yeah i think it would be 13+10*20

Shadow:

Oh, facepalm, 13 is the side of the triangle, not height, LOL

twix01:

okay, thank you!!!

eatsalad:

If not 13+10*20 Then it would be 13+13*20

eatsalad:

yw

Shadow:

You would need the area of the two triangles, two side rectangles, and then bottom rectangle. \[T = (13 \times 10) + (13 \times 20 \times 2) + (10 \times 20)\]